OpenStudy (dillonmerv):

Find the maximum and minimum values of f(x,y)=3x+y on the ellipse x^2+36y^2=1?

OpenStudy (dillonmerv):

I found the value of λ to be ± (5√(13))/12. Is this correct? What do I do next? Can you please show the steps?

ganeshie8 (ganeshie8):

\(3=2x\lambda\) \(1=72y\lambda\) \(x^2+36y^2 = 1\) Dividing first two equations gives \(3 = \dfrac{2x}{72y}\)

ganeshie8 (ganeshie8):

The solutions you've got for \(\lambda\) are correct.

ganeshie8 (ganeshie8):

You may use them to find the \(x,y\) values of the critical points.

OpenStudy (dillonmerv):

So if I equate 3x+y because now I can find the value of x and y through λ, I get ± .0092450033

OpenStudy (dillonmerv):

what is this number? what do I do with this?

ganeshie8 (ganeshie8):

You get two points : x = 0.9985 y = 0.0092 and x = -0.9985 y = -0.0092

ganeshie8 (ganeshie8):

Those two points lie on the given ellipse x^2+36y^2=1 and maximize/minimize the given function f(x,y) = 3x+y.

ganeshie8 (ganeshie8):

Max/min value of 3x+y can only occur at those two points on the ellipse. It cannot occur anywhere else.

OpenStudy (dillonmerv):

Got it Got it!!

OpenStudy (dillonmerv):

I should've known that. Thanks

OpenStudy (dillonmerv):

So I'll just plug those points into 3x+y and I'll get max and min.

OpenStudy (dillonmerv):

The negative should be min